# The Dynamics of Particle-Particle Correlations and the Ridge Effect in Proton-Proton Collisions

### v1

@@ 1 Introduction
Despite the huge theoretical and experimental effort, the ridge effect in proton-proton collisions is still not fully understood [18, 19, 20] despite some recent progress [21, 22, 23]. At present, most of the possible explanations in the literature involve either the Color Glass Condensate (CGC) framework [^1] For a very nice summary, see Ref. [24]. [25, 26, 27] and gluonic flux tubes [28, 29] or hydrodynamic flow [30, 31], while numerous works in the literature offer various descriptions of possible mechanisms responsible for the ridge effect, see for example Refs. [32, 33, 34, 35, 36, 37, 38] and references therein.
Recently, a number of experimental studies has given valuable new information but has not resolved the mystery of the ridge effect in small systems. Collective behaviour of final-state hadrons was studied in high-multiplicity events at photoproduction and deep inelastic $ep$ scattering at a centre-of-mass energy of 318 GeV with the ZEUS detector at HERA [39]. In that study, neither the measurements in photoproduction processes nor those in neutral current deep inelastic scattering showed significant collective behaviour similar to what was observed in high-multiplicity hadronic collisions. Furthermore, measurements of two-particle angular correlations of charged particles emitted in hadronic Z decays were presented in Ref. [40]. The analysis was done with archived $e^{+}$ $e^{-}$ annihilation data at a center-of-mass energy of 91 GeV which were collected with the ALEPH detector at LEP. There, no significant enhancement of long-range correlations was observed. A subsequent analysis on data with a center-of-mass energy of up to 209 GeV identified a long-range near-side excess in the correlation function when calculating particle kinematic variables with respect to the thrust axis [41]. Last year, ALICE released a study of the ridge yield measured in a hadronic system of similar multiplicity to the multiplicity of events that come from $e^{+}$ $e^{-}$ annihilation. The ridge yield was substantially larger than what was observed by the ALEPH analysis for center-of-mass energy of 91 GeV. All these are quite indicative that the mechanisms for ridge yield production in very small hadronic systems are not understood and that more theoretical investigation is needed.
@@
In this paper, we use Monte Carlo simultations to compute the rapidity-azimuthal angle correlations for proton-proton collisions, in particular we use Pythia8[43] as a base reference and BFKLex which is a Monte Carlo code that generates the hard scattering part of the collision using the the Balitsky-Fadin-Kuraev-Lipatov (BFKL) resummation framework [44, 45, 46, 47, 48, 49]. What we aim at with this study is to find out whether any near side correlations arise once we switch from a fixed order $+$ parton shower approach to a BFKL based calculation.
@@ 3 Results and Discussion
Figure 3: Left: Parton angular correlation distribution obtained from p-p collisions with BFKLex. The minimum transverse momentum cutoff of of the outermost in rapidity jets was set to $p_{T,\text{min}}=5$ GeV. The low $p_{T}$ cutoff in fastjet was set to $p^{\text{jet}}_{T,\text{min}}=5$ GeV and the jet radius to $R=0.5$. Right: The same as to the left with the only difference that the low $p_{T}$ cutoff in fastjet was set to $p^{\text{jet}}_{T,\text{min}}=10$ GeV
Figure 4: Left: Parton azimuthal correlation distribution obtained after integrating the distribution in Fig. 3 (Left) over $72<|\Delta\eta|<$ 7. Right: Minijet azimuthal correlation distribution obtained after integrating the distribution in Fig. 3 (Right) over $72<|\Delta\eta|<$ 7.
@@ 4 Conclusion
We presented the first BFKL based Monte Carlo study, directly in momentum space, of the rapidity-azimuthal angle correlations between minijets in proton-proton collisions, together with a similar study with Pythia, the latter serving as a base reference. We found no indication that the particular dynamics of the high-energy limit of QCD can be behind the long-range rapidity correlations and the puzzling ridge effect in small systems. It remains to be seen whether this can change with the inclusion of higher order corrections in a full NLL accuracy BFKL analysis.

### v2

@@ 1 Introduction
Despite the huge theoretical and experimental effort, the ridge effect in proton-proton collisions is still not fully understood [18, 19, 20] despite some recent progress [21, 22, 23]. At present, most of the possible explanations in the literature (for a very nice summary, see Ref. [24]) involve either the Color Glass Condensate (CGC) framework [25, 26, 27] and gluonic flux tubes [28, 29] or hydrodynamic flow [30, 31], while numerous works offer various descriptions of possible mechanisms responsible for the ridge effect, see for example Refs. [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46] and references therein.
Recently, a number of experimental studies have given valuable new information but have not resolved the mystery of the ridge effect in small systems. Collective behaviour of final-state hadrons was studied in high-multiplicity events at photoproduction and deep inelastic $ep$ scattering at a centre-of-mass energy of 318 GeV with the ZEUS detector at HERA [47]. In that study, neither the measurements in photoproduction processes nor those in neutral current deep inelastic scattering showed significant collective behaviour similar to what was observed in high-multiplicity hadronic collisions. Furthermore, measurements of two-particle angular correlations of charged particles emitted in hadronic Z decays were presented in Ref. [48]. The analysis was done with archived $e^{+}$ $e^{-}$ annihilation data at a center-of-mass energy of 91 GeV which were collected with the ALEPH detector at LEP. There, no significant enhancement of long-range correlations was observed. A subsequent analysis on data with a center-of-mass energy of up to 209 GeV identified a long-range near-side excess in the correlation function when calculating particle kinematic variables with respect to the thrust axis [49]. Very recently, ALICE released a study [50] (see also Ref. [51]) of the ridge yield measured in a hadronic system of similar multiplicity to the multiplicity of events that come from $e^{+}$ $e^{-}$ annihilation. The ridge yield was substantially larger than what was observed by the ALEPH analysis for center-of-mass energy of 91 GeV. Furthermore, CMS has reported the observation of enhanced long-range elliptic anisotropies inside high-multiplicity jets in pp collisions [52]. All these are quite indicative that the mechanisms for ridge yield production in very small hadronic systems are not understood and that more theoretical investigation is needed.
@@
In this paper, we use Monte Carlo simultations to compute the rapidity-azimuthal angle correlations for proton-proton collisions, in particular we use Pythia8 [54] as a base reference and BFKLex [55, 56, 57, 58, 59, 60, 61, 62, 63] which is a Monte Carlo code that generates the hard scattering part of the collision using the Balitsky-Fadin-Kuraev-Lipatov (BFKL) resummation framework [64, 65, 66, 67, 68, 69]. In particular, it employs the iterative solution of the BFKL equation cast in a suitable form for Monte Carlo studies directly in momentum space [70]. Both computations are performed within the collinear factorization scheme [71, 72] where the partonic cross section is convoluted with the parton distribution functions (PDF) of the proton. What we aim at with this study is to find out whether any near side correlations arise once we switch from a [LO matrix elements] $+$ [parton shower] approach to a BFKL-based calculation at leading logarithmic accuracy. The importance of computing the correlation distribution from both approaches directly in momentum space from final-state jets and minijets cannot be overstated, as it allows us to avoid any modelling considerations.
@@ 3 Results and Discussion
Figure 3: Left: Minijet angular correlation distribution obtained from p-p collisions with BFKLex. The minimum transverse momentum cutoff of the outermost in rapidity jets was set to $p_{T,\text{min}}=5$ GeV. The low $p_{T}$ cutoff in fastjet was set to $p^{\text{jet}}_{T,\text{min}}=5$ GeV and the jet radius to $R=0.5$. Right: The same as to the left with the only difference that the low $p_{T}$ cutoff in fastjet was set to $p^{\text{jet}}_{T,\text{min}}=10$ GeV.
Figure 4: Left: Minijet azimuthal correlation distribution obtained after integrating the distribution in Fig. 3 (Left) over $72<|\Delta\eta|<$ 7. Right: Minijet azimuthal correlation distribution obtained after integrating the distribution in Fig. 3 (Right) over $72<|\Delta\eta|<$ 7.
@@ 4 Conclusion
We presented the first BFKL based Monte Carlo study, directly in momentum space, of the rapidity-azimuthal angle correlations between minijets in proton-proton collisions, together with a similar study with Pythia, the latter serving as a base reference. We found no indication that the specific dynamics of the high-energy limit of QCD are responsible for the long-range rapidity correlations and the puzzling ridge effect in small systems. It remains to be seen whether this conclusion will change with the inclusion of higher-order corrections in a full NLL accuracy BFKL analysis or with the incorporation of non-linear effects such as parton saturation.